# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ Appendix B Ground Truth Correction with GPSLabeler
Figure 10: Example trajectory fixes with GPSLabeler . The original GPS points are in gray, and the hand-placed markers are blue. The new markers are interpolated with a cubic spline with a color that represents the instantaneous speed: red for slow, and green for fast movement.
In response to this challenge, we have developed the desktop tool GPSLabeler , designed to facilitate the correction of noisy GPS data. The primary strength of GPSLabeler lies in its unique user interface which juxtaposes the video feed and the associated GPS location points on a map. This side-by-side presentation enables users to cross-verify the video content with the GPS coordinates visually.
Using the application is intuitive and straightforward, an example of which can be seen in Figure 3 in the main paper. As users watch the video playback, they can simultaneously observe the GPS path on the map. If discrepancies are noted, users can manually re-label the GPS data by simply clicking on the map and dragging the points to the desired accurate location. One key feature of GPSLabeler is its interpolation mechanism. Once users adjust certain key points on the map, the application automatically interpolates the GPS data between these points, ensuring smooth and consistent transitions.
Designed with user-friendliness in mind, GPSLabeler is cross-platform, with versions available for Windows, MacOS, and Linux. This ensures that researchers and dataset curators across diverse technological environments can access and utilize the tool to enhance the accuracy of their geospatial data.
@@
Appendix C Dataset Postprocessing
C.1 Gaze Fixation Detection
For our fixation detection, we adopted the methodology from Pupil Labs, the creators of the glasses embedded with gaze detectors [43]. Our criteria define gaze fixations as instances where the gaze remains relatively stationary for a duration ranging from 80ms to 1 second. The upper limit of this range, 1 second, could also characterize a smooth pursuit, where the eyes smoothly track a moving object. Furthermore, we account for dispersion, which represents the spatial movement permitted within a single fixation. A gaze group is considered a fixation if the dispersion is within a limit of 1.5 degrees, ensuring that minor involuntary or random gaze movements do not undermine the detection of genuine fixations.
C.2 Exploring Image Homography and Stitching
  • Using the current state of the art, it is not feasible to register the driver head-cam frames and the scene videos within an acceptable error margin. We experimented with classical feature extractors such as DISK [44], as well as transformer-based detector-free local feature matching [41], using Kornia library [33].
  • We have observed that the redundancy from having multiple scene camera feeds can be exploited for training regularization. By dropping out one of the views at each frame during training, we were able to improve the validation performance.
Example frames from the stitching experiments can be seen in Figure 11.
Scene 1
in 0,10,…,90
Scene 2
in 0,10,…,90
Scene 3
in 0,10,…,90
Scene 4
in 0,10,…,90
Scene 5
in 0,10,…,90
(a) Frame 1
Figure 13: Two simple intersections where both gaze and video models are successful. In example (a), there is a 90-degree right turn with the correct curvature estimated by both the gaze and video models. G-MEMP is better at predicting the speed for the later stages of the trajectory. In example (b), both models have a near-identical prediction for a sharp merge from a residential road to a main street.
Appendix D G-MEMP
$\Psi\left(M_{1:T},S_{1:T}^{1:K},G_{1:T}|\lambda\right)\equiv$
$\mathcal{T}\left(E_{M}(M_{1:T}),E_{P}(E_{S}(S_{1:T}^{1:K}),E_{G}(G_{1:T}))\right)=$
$\qquad\qquad\qquad\qquad\qquad\{M^{\prime}_{T:T_{\text{pred}}},s_{T:T_{\text{pred}}}^{1:K},g_{T:T_{\text{pred}}}\}$
Loss definitions.
For the model outputs $\Psi\left(M_{1:T},S_{1:T}^{1:K},G_{1:T}|\lambda\right)=\{M^{\prime}_{T:T_{\text{pred}}},s_{T:T_{\text{pred}}}^{1:K},g_{T:T_{\text{pred}}}\}$, let $\mathcal{L}_{T}$ denote the discounted trajectory loss, $\mathcal{L}_{V}$ the discounted video loss and $\mathcal{L}_{G}$ the discounted gaze loss. The auxiliary losses are defined as:
$\mathcal{L}_{V}$ $=\sum_{i=1}^{K}\mathcal{L}_{\text{fd}}\left(s^{i}_{T:T_{\text{pred}}},E_{S}(S_{T:T_{\text{pred}}}^{i})\right)$
$\mathcal{L}_{G}$ $=\mathcal{L}_{\text{fd}}\left(g_{T:T_{\text{pred}}},E_{G}(G_{T:T_{\text{pred}}})\right).$
D.1 Implementation Details
D.2 Module-level Caching with ModuleCache
To make this process repeatable, we implemented the ModuleCache utility to provide a generic caching mechanism. The ModuleCache library offers an easy way to cache the outputs of PyTorch modules, making it particularly suited for our need to efficiently cache the outputs of large, pre-trained vision transformers.
D.2.1 Key Features
The ModuleCache library offers several notable features:
D.2.2 Usage
The library’s decorator-based design ensures minimal modifications to the existing PyTorch codebase. By adding the @modulecache() decorator, the output of the module’s forward method gets cached, as illustrated in the provided basic usage example:
1from @
2class MyModule(nn.Module):
3    def init(self):
4        super().init()
5        self.linear = nn.Linear(10, 10)
6    def forward(self, x):
7        # This output will be cached
8        return self.linear(x)
More detailed examples, advanced usage patterns, and further documentation will be accessible at the official ModuleCache documentation.
D.2.3 Operational Assumptions
For the ModuleCache library to operate seamlessly, there are certain assumptions that the PyTorch modules should adhere to:
  1. 1.
    The module should be a subclass of nn.Module.
  2. 2.
    The forward method of the module should be capable of accepting any number of positional arguments with shapes denoted by $(B,*)$, where $B$ stands for the batch size, and *** denotes any other dimensions.
  3. 3.
  4. 4.
    The method should return a single tensor with the shape $(B,*)$.
The ModuleCache library has proven invaluable in our experiments by allowing us to efficiently utilize the outputs of our pre-trained vision transformers without incurring significant computational overhead.
Appendix E Experiments
E.1 Evaluation Metrics
The Average Displacement Error (ADE) is formally defined as:
E.2 Hyperparameters
Table 6: Training configurations
Parameter Value
Optimizer AdamW
Learning rate 0.0005
Weight decay 0.05
Epochs 250
Epsilon 1
Scene video FPS 1
Gaze video FPS 1
Batch size 4
Output FPS 5
Min PCI 20
Loss discount factor/epoch [0: 0.9, 100: 0.95, 200: 0.99]
Table 7: Framework configurations
Parameter Value
Encoder heads 8
Encoder layers 3
Feature dropout 0.1
Dense prediction true
Use angle and norm true
Encoder hidden size 64
Gaze encoder layers 3
Gaze dropout 0.5
View dropout 1
Maxpool kernel size 8
Front video scaling factor 0.3
Scene video scaling factor 0.1
Gaze decoder dropout 0.1
Image embedding size 64
Dense gaze loss ratio 0.2
Dense video loss ratio 0.2
Normalize video features true
Table 8: GPS backbone (Informer) configurations
Parameter Value
Distil true
Factor 1
Model dimension 128
Dropout 0.1
Number of Heads 8
Sequence Length 40
Decoder layers 2
Encoder layers 3
Prediction Length 30
Label Length 40
With Gaze true
Activation gelu
Individual false
Moving average 25
Table 9: Video backbone (SwinV2) configuration
Parameter Value
Model Type swinv2_base_window12to16_192to256
Pad to square true
Train backbone false
E.3 Qualitative analysis of MEMP
We have shown that by utilizing gaze, G-MEMP is able to make better predictions in certain (high path complexity) situations, as explored in Figure 7. While driver intention prediction is important in these cases, we notice that for many simpler straightforward cases, predictions based on scene and trajectory information are already quite accurate. In Figure 13, we demonstrate the performance of MEMP. In these examples, there is an obvious turn that the driver can make, and the curvature of the trajectory is determined by the roads in front, which MEMP model is successful in figuring out without extra driver information.

### v2

@@ C.1 Ground Truth Correction with TrajApp
Figure 10: Example trajectory fixes with TrajApp . The original GPS points are in gray, and the hand-placed markers are blue. The new markers are interpolated with a cubic spline with a color that represents the instantaneous speed: red for slow, and green for fast movement.
Figure 12: GPS distances to the nearest road center. Estimated via OpenStreetMap. DR(eye)VE’s GPS points are further away from road centers, indicating significant noise.
C.2 GEM vs. DR(eye)VE Noise Comparison
To demonstrate the difference in degree of noise between the proposed GEM dataset and the driving gaze saliency dataset from the literature, DR(eye)VE, we plotted the distribution of the GPS distances to the nearest road center for each dataset in Figure 12. The road centers are gathered from the public data in OpenStreetMap. For GEM, the distance distribution is centered around 1.25 meters, which is consistent with the average lane width of 3 meters in urban settings, while the distribution is considerably more heavy-tailed in the case of DR(eye)VE, indicating significant GPS noise.
C.3 Gaze Fixation Detection
For our fixation detection, we adopted the methodology from Pupil Labs, the creators of the glasses embedded with gaze detectors (Tonsen et al., 2020). Our criteria define gaze fixations as instances where the gaze remains relatively stationary for a duration ranging from 80ms to 1 second. The upper limit of this range, 1 second, could also characterize a smooth pursuit, where the eyes smoothly track a moving object. Furthermore, we account for dispersion, which represents the spatial movement permitted within a single fixation. A gaze group is considered a fixation if the dispersion is within a limit of 1.5 degrees, ensuring that minor involuntary or random gaze movements do not undermine the detection of genuine fixations.
C.4 Gaze Alignment and Robustness
Our framework processes driver gaze data following established practices in eye-tracking research (Hayhoe et al., 2003; Land, 2006). To align the high-frequency gaze measurements (200 Hz) with video frames (5 Hz), we use median downsampling. This approach ensures robust position estimates while preserving the temporal characteristics of driver attention patterns.
To evaluate our model’s robustness to potential measurement errors, we conducted additional experiments on the GEM dataset. As shown in Table 12, artificially introducing gaze position noise of $\pm$50 pixels (approximately 5% error) results in minimal degradation of prediction accuracy (0.01m increase in ADE). This resilience can be attributed to our "field-of-view" approach, which treats gaze as supplementary to the driver’s head-mounted video feed rather than as a primary input.
Configuration ADE (m) ADE+20PCI (m)
RouteFormer - without noise 5.99 7.70
RouteFormer - with $\pm$50px noise 6.00 7.72
RouteFormer-Base (without FOV) 6.13 8.14
Table 12: Impact of gaze measurement noise on prediction accuracy. The minimal change in ADE demonstrates our model’s robustness to gaze uncertainty. Top and bottom rows are from Table 2
These results demonstrate that our processing pipeline, built on standard gaze analysis techniques, effectively handles both natural gaze patterns and potential measurement uncertainties.
C.5 Exploring Image Homography and Stitching
  • Using the current state of the art, it is not feasible to register the driver headcam and the scene videos within an acceptable error margin. We experimented with classical feature extractors such as DISK (Tyszkiewicz et al., 2020) and transformer-based detector-free local feature matching (Sun et al., 2021), using Kornia library (Riba et al., 2019).
  • We have observed that the redundancy from having multiple scene camera feeds can be exploited for training regularization. By dropping out one of the views at each frame during training, we improved the validation performance.
Example frames from the stitching experiments can be seen in Figure 13.
Scene 1
[…]
in 0,10,…,90
Scene 2
[…]
in 0,10,…,90
Scene 3
[…]
in 0,10,…,90
Scene 4
[…]
in 0,10,…,90
Scene 5
[…]
in 0,10,…,90
(a) Frame 1
Appendix D PCI Metric
D.1 Breakdown of GEM Dataset and PCI Insights
A deeper dive into our dataset, combined with synchronization with OpenStreetMap (OpenStreetMap contributors, 2017), reveals that the high PCI events indeed happen either close to the intersections, or on trajectories with a high-speed change, as observed in Figure 15.
@@ D.1 Breakdown of GEM Dataset and PCI Insights
D.2 Visualizing Path Complexity Index
Figure 16 illustrates the key components of PCI computation. Given an input trajectory (gray), we extrapolate a simple trajectory (dashed) by maintaining the final velocity vector. PCI measures the Fréchet distance (red) in meters between this simple trajectory and the actual target path (blue), effectively quantifying how much the actual path deviates from the simple projection. This deviation captures the complexity of the maneuver - higher deviations indicate more complex trajectories, requiring additional context to predict accurately. Unlike an alternative metric like mean-squared error, Fréchet distance is more stable against speed changes, better capturing differences in trajectory rather than point-by-point error.
Figure 16: A simple representation of PCI as a concept.
D.3 PCI Analysis on Urban Driving
Figure 17 visualizes PCI values across a representative urban drive in GEM. High PCI regions (bright yellow) consistently align with challenging scenarios: roundabout exits (top-right, PCI $>$ 80), lane merges after intersections (bottom-left, PCI $>$ 60), and sharp turns exceeding 45 degrees (top-left and bottom-right). Notably, slower 90-degree turns exhibit lower PCI values due to reduced divergence from simple trajectories, demonstrating PCI’s ability to distinguish the degree of challenge to predict each target curve.
Figure 17: Average PCI values of a full trajectory of a video in GEM. Estimated by running a sliding window of 1 second across the whole trajectory with 6s-8s input-target length, and averaging the PCI values assigned to each point. Brightest yellow parts map to $\sim 80$ PCI while the dark blue is at 0 PCI.